# Task-optimal data-driven surrogate models for eNMPC via differentiable simulation and optimization

### v1

@@ III-B Training setup
    Koopman-SI: eNMPC controller using a Koopman surrogate model trained using SI.
@@
    Koopman-PPO: eNMPC controller using a Koopman surrogate model pretrained using SI and refined for task-optimal performance using the state-of-the-art policy gradient algorithm Proximal Policy Optimization (PPO) [24].
@@
    Koopman-SHAC (main contribution of this work): eNMPC controller using a Koopman surrogate model pretrained using SI and refined for task-optimal performance using the SHAC algorithm [7].
  4. 4.
    MLP-PPO: Neural network controller in the form of a multilayer perceptron (MLP) trained using PPO.
  5. 5.
    MLP-SHAC: MLP controller trained using SHAC.
A common practice when training prediction models is to oversize the models and prevent overfitting through regularization techniques and early stopping of the training. However, our ultimate goal is to train task-optimal computationally cheap dynamic surrogate models. Therefore, we avoid oversizing of the Koopman models by starting with a small model and iteratively training bigger models until the gains in prediction accuracy in the SI pretraining become negligible. We take an analogous approach to identify a suitable prediction horizon of the eNMPC controller. For brevity, we waive showing the results of these preprocessing steps. All results presented in Subsection III-C are obtained using an eNMPC horizon of nine hours and a Koopman model with a latent space dimensionality of eight, i.e., $\bm{A}_{\bm{\theta}}\in\mathbb{R}^{8\times 8},\bm{B}_{\bm{\theta}}\in\mathbb{R}^{8\times 2},\bm{C}_{\bm{\theta}}\in\mathbb{R}^{2\times 8}$, and an MLP encoder $\bm{\psi}\colon\mathbb{R}^{2}\mapsto\mathbb{R}^{8}$ (two hidden layers, four and six neurons, respectively; hyperbolic tangent activation functions). The MLP controllers have three distinct input layers: one for the states of the CSTR model, one for the storage level, and one for the trajectory of future electricity prices. We concatenate the outputs of those layers and pass them through two hidden layers of size 64 neurons. The output layer has two neurons, one each for the control variables $\rho$ and $F$.
We use the same data set as in [4] for the SI pretraining of the Koopman surrogate model. This data set consists of 84 trajectories, each of a length of 5 days, i.e., 480 time steps, using a step length of 15 minutes. Of those 84 trajectories, we use 63 for training and the remaining 21 for validation. We perform SI by minimizing the sum of the three loss terms presented in Subsection II-B (Eq. 3-5). We repeat SI ten times using random seeds. We use the model with the lowest validation loss for the Koopman-SI controller. The same model is used as the initial guess when training the Koopman-PPO and Koopman-SHAC controllers.
In order to use a policy optimization algorithm such as SHAC, which makes use of derivative information from a simulated environment, not only the dynamic model but also the reward function must be differentiable. For our case study, we choose a reward function that calculates the reward at time step $t$ based on whether any constraints were violated at that time step, and on the electricity cost savings compared to steady-state production between $t-1$ and $t$. The constraint component $r^{\text{con}}_{t}$ of the reward quadratically penalizes violations of the bounds of $c$, $T$, and the product storage, i.e., $r^{\text{con}}_{t}\geq 0$, and $r^{\text{con}}_{t}=0$ if no constraint violation occurs at $t$. The cost-component $r^{\text{cost}}_{t}$ of the reward is calculated by taking the difference between the cost at steady-state production and the actual cost between $t-1$ and $t$:
@@
$r_{t}=\alpha\cdot r^{\text{cost}}_{t}-r^{\text{con}}_{t}.$
Figure 3: Comparison of the controllers. We show a randomly chosen two-day interval from the test episode. The dotted black lines represent the bounds.
@@
We train the policies using day-ahead electricity prices from the Austrian market from March 29, 2015, to March 25, 2018 [25]. Using random seeds, we repeat the controller training ten times for each combination of policy and training algorithm (except for the Koopman-SI controller, whose Koopman model is not trained any further after SI).
All code used for training the controllers, including the hyperparameters that were used to obtain the results presented in Subsection III-C, is publicly available on GitLab [^1] https://jugit.fz-juelich.de/iek-10/public/optimization/shac4koopmanenmpc. We used the Stable-Baselines3 [26] implementation of PPO but implemented our own version of SHAC following the description in [7].
III-C Results
@@
For each type of controller trained using PPO or SHAC, we identify the controller (and the associated set of parameters) that achieved the highest average reward between two consecutive parameter updates. We test their performance and that of the Koopman-SI controller on a continuous roughly half-year-long test episode using Austrian day ahead electricity price data from March 26, 2018, to September 30, 2018 [25]. Unlike the training process, we perform this testing without exploration, i.e., we waive adding Gaussian noise to the controller output (cf. Section II-D). We initialize the test episode for each controller at the steady state of the CSTR and with empty product storage. The results are presented in Table I. We also visualize the behavior of each controller by randomly choosing a two-day interval from the test episode and plotting the product concentration $c$, the production rate $\rho$, and the coolant flow rate $F$ during this interval in Fig. 3. It can be noted that all controllers leverage the entire feasible range of the control variables and the product concentration, and that the results exhibit an intuitive inverse relationship between electricity prices and coolant flow rate. We forego including the temperature $T$ in Fig. 3 as $T$ never reaches its bounds in any of the test episodes.
@@
TABLE I: Test results: Costs are stated relative to steady-state production cost. The percentage of control steps that result in constraint violations is given. The average size of a constraint violation is given relative to the size of the feasible range of the variable whose bound was violated.
Cost Constr. viols. [%] Avg. size constr. viol.
MLP-PPO $\bm{0.88}$ $6.34$ $9.3\cdot 10^{-3}$
MLP-SHAC $0.90$ $0.93$ $9.8\cdot 10^{-3}$
Koopman-SI $0.90$ $36.24$ $8.2\cdot 10^{-3}$
Koopman-PPO $0.90$ $22.31$ $9.9\cdot 10^{-3}$
Koopman-SHAC $0.90$ $\bm{0.39}$ $\bm{6.0\cdot 10^{-5}}$
As can be seen from Table I, Koopman-SHAC causes the fewest and also the smallest constraint violations of all tested controllers in this particular case study. MLP-PPO achieves slightly higher cost savings, but produces a much higher rate and size of constraint violations.
@@
Figure 4: Learning progress in the Koopman-SHAC training runs. The dark blue line indicates the running average reward over the previous 1024 steps in the environment, averaged over all ten training runs. The light blue region indicates one standard deviation of the performance variance between the training runs.
@@
It is noticeable that Koopman-SI and Koopman-PPO cause substantially more constraint violations than the other controllers. In the case of Koopman-SI, this is not surprising since the employed Koopman surrogate model received no end-to-end training for task-optimal performance. Here, frequent but minor constraint violations show that the controller is trying to operate at the borders of the feasible range, which is normal behavior for a predictive controller. However, in the case of Koopman-PPO the results are somewhat unsatisfactory and we observe that some training runs did not improve performance compared to Koopman-SI. Moreover, those training runs that did improve performance did not show stable convergence to high rewards. In contrast, SHAC leads to a relatively stable convergence to high rewards in our particular case study (Fig. 4). The control performance improves relatively evenly in all ten training runs without ever dropping for extended periods.
@@
Due to the size of the case study under consideration, a detailed analysis of the training runtimes would provide little insight about the expected runtime on a control problem of more practical relevance. Therefore, we leave such an analysis for future work on larger systems.
IV Conclusion
@@
We combine our previously published method for turning Koopman-(e)NMPC controllers into differentiable policies [4] with the policy optimization algorithm SHAC [7]. SHAC leverages derivative information from automatically differentiable simulated environments. We show that exploiting such derivative information translates to superior performance in end-to-end training of a Koopman surrogate model for eNMPC of a literature known CSTR. We also show that convergence to high control performance is fairly stable across all independent training instances. The results can be understood as a successful proof of concept. Future work should investigate the application of our method to larger mechanistic simulation models and more challenging control problems, presumably necessitating training for more iterations. Furthermore, the computational burden of backpropagation through mechanistic simulations and optimal control problems and thus the cost of each iteration is naturally linked to the size of the mechanistic simulation model, meaning that our method could become computationally intractable for very large models.
ACKNOWLEDGMENT
@@
This work was performed as part of the Helmholtz School for Data Science in Life, Earth and Energy (HDS-LEE) and received funding from the Helmholtz Association of German Research Centres.

### v3

@@ 3.2 Training setup
    Koopman-SI: eNMPC controller using a Koopman surrogate model trained solely using SI.
@@
    Koopman-PPO: eNMPC controller using a Koopman surrogate model pretrained using SI and refined for task-optimal performance using the state-of-the-art policy gradient algorithm Proximal Policy Optimization (PPO) (Schulman et al., (2017)), like we did in Mayfrank et al., (2024).
@@
    Koopman-SHAC (main contribution of this work): eNMPC controller using a Koopman surrogate model pretrained using SI and refined for task-optimal performance using the SHAC algorithm (Xu et al., (2022)).
Our goal is to train dynamic surrogate models of fixed size for optimal performance as part of eNMPC. All results presented in Subsection 3.3 are obtained using a Koopman model with a latent space dimensionality of eight, i.e., $\bm{A}_{\bm{\theta}}\in\mathbb{R}^{8\times 8},\bm{B}_{\bm{\theta}}\in\mathbb{R}^{8\times 2},\bm{C}_{\bm{\theta}}\in\mathbb{R}^{2\times 8}$, and an encoder $\bm{\psi}\colon\mathbb{R}^{2}\to\mathbb{R}^{8}$ in the form of a feedforward neural network with two hidden layers, four and six neurons, respectively and hyperbolic tangent activation functions. We use an eNMPC horizon of nine hours.
We use the same data set as in (Mayfrank et al., (2024)) for the SI pretraining of the Koopman surrogate model. This data set consists of 84 trajectories, each of a length of 5 days, i.e., 480 time steps, using a step length of 15 minutes. Of those 84 trajectories, we use 63 for training and the remaining 21 for validation. Then, we perform SI of the Koopman model in a similar way as described in (Mayfrank et al., (2024)). We repeat SI ten times using random seeds. We use the model with the lowest validation loss for the Koopman-SI controller. The same model is used as the initial guess when training the Koopman-PPO and Koopman-SHAC controllers.
In order to use a policy optimization algorithm such as SHAC, which makes use of derivative information from a simulated environment, not only the dynamic model but also the reward function must be differentiable. For our case study, we choose a reward function that calculates the reward at time step $t$ based on whether any constraints were violated at that time step, and on the electricity cost savings compared to the steady-state production at nominal rate between $t-1$ and $t$. The constraint component $r^{\text{con}}_{t}$ of the reward quadratically penalizes violations of the bounds of $c$, $T$, and the product storage, i.e., $r^{\text{con}}_{t}\geq 0$, and $r^{\text{con}}_{t}=0$ if no constraint violation occurs at $t$. The cost-component $r^{\text{cost}}_{t}$ of the reward is calculated by taking the difference between the cost at nominal production and the actual production cost between $t-1$ and $t$, i.e.,
@@
$r_{t}=\alpha\cdot r^{\text{cost}}_{t}-r^{\text{con}}_{t}$
@@
We train the policies using day-ahead electricity prices from the Austrian market from March 29, 2015, to March 25, 2018 (Open Power System Data, (2020)). Using random seeds, we repeat the controller training ten times for each combination of policy and training algorithm (except for the Koopman-SI controller, whose Koopman model is not trained any further after SI).
We use the same hyperparameters for Koopman-PPO and Koopman-SHAC wherever possible, i.e., for all hyperparameters which are not part of PPO or SHAC. For the hyperparameters of PPO and SHAC, we do not perform extensive hyperparameter tuning. Instead, we mostly rely on the standard values. In both algorithms, we use the Adam optimizer with a small learning rate of $10^{-5}$ as the behavior of the Koopman eNMPCs is highly sensitive to small changes in the parameters. We used the Stable-Baselines3 (Raffin et al., (2021)) implementation of PPO but implemented our own version of SHAC following the description in (Xu et al., (2022)). All code used for training the controllers, including the hyperparameters that were used to obtain the results presented in Subsection 3.3, is publicly available [^1] https://jugit.fz-juelich.de/iek-10/public/optimization/shac4koopmanenmpc. In addition to the code used to obtain the results presented in the following subsection, the linked repository also contains code for training pure neural network policies for our case study using PPO and SHAC. As the results of the neural network policies do not influence the narrative of this work in a meaningful way we do not discuss these results.
3.3 Results
@@
For Koopman-PPO and Koopman-SHAC, we identify the controller (and the associated set of parameters) that achieved the highest average reward between two consecutive parameter updates. We test their performance and that of the Koopman-SI controller on a continuous roughly half-year-long test episode using Austrian day ahead electricity price data from March 26, 2018, to September 30, 2018 (Open Power System Data, (2020)). Unlike the training process, we perform this testing without exploration, i.e., we waive adding Gaussian noise to the controller output (cf. Fig. 3). We initialize the test episode for each controller at the steady state of the CSTR and with empty product storage. The results are presented in Table 1. The trajectories produced by all controllers exhibit an intuitive inverse relationship between electricity prices and coolant flow rate.
@@
Table 1: Test results: The economic cost is stated relative to the nominal production cost, i.e., we report the cost incurred by the respective controller, divided by the cost incurred by steady-state production at nominal rate given the same electricity price trajectory. The percentage of control steps that result in constraint violations is given. The average size of a constraint violation is given relative to the size of the feasible range of the variable whose bound was violated.
Economic cost Constr. viols. [%] Avg. size constr. viol.
Koopman-SI $\bm{0.88}$ $19.88$ $5.1\cdot 10^{-2}$
Koopman-PPO $0.90$ $17.57$ $1.3\cdot 10^{-2}$
Koopman-SHAC $0.90$ $\bm{0.0}$${-}\qquad$
As can be seen from Table 1, Koopman-SHAC is the only controller that does not produce any constraint violations. Koopman-SI achieves slightly higher cost savings than Koopman-SHAC, however, it frequently produces constraint violations. Therefore, we consider Koopman-SHAC preferable in most real-world applications where constraint-satisfaction is of high importance.
@@
Figure 4: Learning progress in the Koopman-SHAC training runs. The dark orange line indicates the running average reward over the previous 1024 steps in the environment, averaged over all ten training runs. The light orange region indicates one standard deviation of the performance variance between the training runs.
@@
It is noticeable that Koopman-SI and Koopman-PPO cause substantially more constraint violations than the Koopman-SHAC controller. In the case of Koopman-SI, this is not surprising since the employed Koopman surrogate model received no end-to-end training for task-optimal performance. Here, frequent but minor constraint violations show that the controller is trying to operate at the borders of the feasible range, which is normal behavior for a predictive controller. However, in the case of Koopman-PPO the results are unsatisfactory and rather unexpected. Here, the end-to-end training reduced the average size of the constraint violations by roughly a factor of four, but only led to a small improvement in the frequency of constraint violations. We observe that some training runs did not improve performance compared to Koopman-SI at all. Moreover, those training runs that did improve performance did not show stable convergence to high rewards. The unstable convergence of the Koopman-PPO controllers is in line with the results in our previous work (Mayfrank et al., (2024)). There, we used a non-differentiable reward formulation with a high constant penalty incurred by any constraint violation. Using that reward formulation, we managed to substantially reduce the frequency of constraint violations unlike here. However, that approach (Mayfrank et al., (2024)) severely affected the resulting economic performance and thus produced overly conservative eNMPC controllers. In contrast to Koopman-PPO, Koopman-SHAC exhibits a relatively stable convergence to high rewards in our case study (Fig. 4) and produces superior terminal performance (Tab. 1). The control performance improves relatively evenly in all ten training runs without ever dropping for extended periods.
@@
Due to the small size of the case study under consideration, a detailed analysis of the training runtimes would provide little insight about the expected runtime on a control problem of more practical relevance. Therefore, we leave such an analysis for future work on larger systems.
3.3.1 Policy gradient analysis
The fundamental difference between the two policy optimization algorithms SHAC (Xu et al., (2022)) and PPO (Schulman et al., (2017)) is that SHAC utilizes analytic policy gradients from an automatically differentiable environment, whereas PPO estimates policy gradients via the policy gradient theorem (Sutton and Barto, (2018)). Ilyas et al., (2018) show that given common and practical hyperparameter configurations, the PPO-estimated policy gradients can incur a high variance which can lead to unstable training convergence, as we observe in our case. We follow the approach by Ilyas et al., (2018) to analyze empirically whether there is a meaningful difference in the variance of the policy gradients produced by PPO and SHAC when applied to our case study. To this end, we investigate how similar the direction of multiple policy gradients are given a fixed policy parameterization. We fix the policy parameters to that of the Koopman-SI controller for this analysis. Then, for PPO and SHAC, we fit the critic to this policy without updating the policy. Finally, for both algorithms and still without updating the policy, we record the policy gradients of 100 optimization steps. We calculate how similar the 100 recorded gradients are by computing their average pairwise cosine similarity. The cosine similarity is a measure of the similarity of two vectors which only depends on their direction, not on their length. It takes a value of one if both vectors point in the same direction, zero if they are orthogonal to each other, and minus one if they point in exactly opposite directions. PPO produces an average cosine similarity of $0.22$, whereas SHAC results in an average cosine similarity of $0.94$. Thus, the variance in the direction of policy gradients is much higher for PPO than for SHAC, which might explain the instable convergence to high rewards and thus the relatively bad performance of Koopman-PPO (see Tab. 1).
4 Conclusion
We combine our previously published method for turning Koopman-(e)NMPC controllers into differentiable policies (Mayfrank et al., (2024)) with the policy optimization algorithm SHAC (Xu et al., (2022)). Our approach leverages derivative information from automatically differentiable simulated environments, differentiating it from previously published methods for end-to-end training of dynamic models for control. We find that SHAC produces a stable convergence to high control performance across all independent training instances, translating to superior control performance compared to our previously published approach (Mayfrank et al., (2024)) which was based on the PPO algorithm. Note that even though our method achieves perfect constraint satisfaction in our case study, full adherence to constraints can in general not be expected.
@@ 4 Conclusion
The results can be understood as a successful proof of concept. By training data-driven surrogate models for optimal performance in a specific control task, our method utilizes the representational capacity of the model efficiently, thus avoiding unnecessarily large and computationally expensive surrogate models. We view our approach as a promising avenue toward more performant real-time capable data-driven (e)NMPCs. Still, the computational burden of backpropagation through mechanistic simulations and optimal control problems and thus the cost of each training iteration is naturally linked to the size of the mechanistic simulation model, meaning that our method could become computationally intractable for very large models. Thus, future work should investigate the application of our method to larger mechanistic simulation models and more challenging control problems, presumably necessitating training for more iterations.
Declaration of Competing Interest
We have no conflict of interest.
Acknowledgements
@@
This work was performed as part of the Helmholtz School for Data Science in Life, Earth and Energy (HDS-LEE) and received funding from the Helmholtz Association of German Research Centers.